# Nordlund0000-0001-6244-1942

### v1

@@ VI.2 Comparison of range profiles with experiments
Root-mean-square deviation 4.5% 5.5% 6.1% 5.0% 6.8%
@@ VII Summary and conclusions
We have discussed calculations of repulsive interatomic potentials at three levels of theory. The one-electron basis set is known to be the most important source of error for repulsive potentials [43]. A numerically robust basis set scheme has been recently verified against fully numerical electronic structure calculations at the Hartree–Fock level of theory in Ref. 41. The state of the art was represented in this work by an analogous scheme based on the use of flexible Gaussian-type orbital (GTO) basis sets within the PySCF program [51], in which the total energy was estimated with second-order Møller–Plesset perturbation theory (MP2) calculations. The MP2 calculations were carried out for the set of diatomic molecules with $Z_{1}+Z_{2}\leq 36$.
@@
The third level of theory considered was orbital-free density functional theory (OF-DFT) evaluated on the superposition of atomic densities, following the approach of Ziegler, Biersack, and Littmark (ZBL) in Ref. 13. We found that the ZBL potentials and the reproduced OF-DFT calculations have considerable differences to our more accurate self-consistent quantum chemical calculations: both the ZBL universal potential and the ZBL pair-specific calculations differ from the quantum chemical approaches by $\sim$ 5%.
Having verified the accuracy of the DMol scheme, we formed pair-specific analytical fits of the DMol screening functions to the sum of three exponentials, yielding the purely repulsive NLH potentials of this work. By comparison to the MP2 reference data, we showed that the pair-specific NLH potentials agree with the best available quantum chemical data within $\sim 2$% above 30 eV. The exponents and fitting coefficients defining the NLH potential are included in the Supplementary Material and in an Open data set.
@@
The ZBL pair-specific, DMol and MP2 data sets and the NLH potential coefficients are available open access in the online Supplementary materials of this publication in the compressed archive package file nlh_potentials_opendata.tar.gz , as well as at https://zenodo.org/records/14172633 (doi:0.5281/zenodo.14172632).
The online supplementary material also includes the file nlh_table_all.pdf that has a table of all the NLH potential coefficients.
@@ Appendix A Combining high-energy and equilibrium potentials
We now address the question of how the two types of interactions—near-equilibrium many-body and high-energy repulsive—can be combined. The key to this combination is that at the highest energies the atom collisions can practically always be handled as two-body interactions, as strong keV energy collisions involve internuclear distances $<0.5$, while regular internuclear distances are of the order of 2–3 Å. It is thus practically impossible for more than two atoms to simultaneously experience the strongly repulsive part of the potential. It is also fully sufficient to use BCA type repulsive pair potentials of the form of eq. 4 for small internuclear distances in MD simulations.
@@
where $0\leq F(r)\leq 1$ is some continuous function that has the properties $F(r)\to 1$ when $r\to 0$ and $F(r)\to 0$ when $r\to\infty$. To achieve a sharp transition from the many-body to the repulsive potential, the function $F(r)$ should transition from 1 to 0 in a narrow interval below $r_{\rm equi}$; the Fermi function is one possibility [99]. This approach has the advantage that it naturally ensures that e.g. angular dependencies associated with chemical bonds are not active at small interatomic separations.
@@ Appendix B Computation of the orbital-free density functional potentials
The basic input to the calculations are the electron densities $\rho(r)$, where $\rho$ is given in units of electrons per Å$\rm\AA^{3}$ and the distance from the nucleus $r$ is in units of $\rm\AA$. We used the spherically averaged solid-state electron densities as listed in the ZBL book [13] to determine what we call the pair-specific ZBL potentials for all non-radioactive elements, plus Bi and U. In addition, we calculated electron densities for many atoms in the gas phase with the Grasp2018 program [44, 45] for comparison.
@@
Equations 23 and 24 describe the interaction of a charge cloud with a point charge, integrating over all of space, while eq. 25 describes the interaction of two charge clouds with each other. In eqs. 23 and 24, $r_{1}$ and $r_{2}$ denote the distances of the infinitesimal volume element ${\rm d}^{3}r$ from nuclei 1 and 2, respectively. Equation 25 is for now written in a general form where ${\bf x}_{1}$ and ${\bf x}_{2}$ are coordinates of two volume elements, and $|{\bf x}_{1}-{\bf x}_{2}|$ is the distance between them. The numerical value of $e^{2}/(4\pi\epsilon_{0})$ was given in the main text after eq. 1.
@@
The Coulomb integrals of eqs. 23, 24 and 25 are facile to evaluate using Green’s theorem or the Laplace expansion for the potential [49]: a spherical charge distribution yields a potential which coincides with that of a point charge
@@
Since we approximate the electron densities $\rho_{i}(r)$ by piecewise polynomials, the integrals in eqs. 26 and 39 can be carried out exactly, which improves the accuracy of the Coulomb interaction term eq. 22. This also speeds up the computation of $V_{c}$, because the integrals can be represented by another piecewise polynomial, which has to be defined only once for each atom species.
The integrals in eqs. 38, 40 and 41 were evaluated using the quad function of the SciPy library [107], using a relative tolerance of $10^{-6}$ for the quantum-mechanical terms and $10^{-7}$ for the Coulomb term $V_{ee}$. For the quantum-mechanical terms, the tolerance cannot always be reached, but the relative error usually is better than $10^{-4}$. In addition, the error estimate is not guaranteed to be accurate. Numerical problems are detected and fixed in the course of choosing the abscissae for the function $V(R)$ as described in the following.

### v3

@@ VI.2.3 Comparison of results with experiments
Root-mean-square deviation 4.3% 5.3% 6.0% 4.9% 6.6%
@@ VII Summary and conclusions
We have presented and compared results of repulsive interatomic potentialcalculations at three levels of theory. The one-electron basis set is known to be the most important source of error for repulsive potentials [43]. A numerically robust basis set scheme has been recently verified against fully numerical electronic structure calculations at the Hartree–Fock level of theory in Ref. 37. The state of the art was represented in this work by an analogous scheme based on the use of flexible Gaussian-type orbital (GTO) basis sets within the PySCF program [51], in which the total energy was estimated with second-order Møller–Plesset perturbation theory (MP2) calculations. The MP2 calculations were carried out for the set of diatomic molecules with $Z_{1}+Z_{2}\leq 36$.
@@
The third level of theory considered was orbital-free density functional theory (OF-DFT) evaluated on the superposition of atomic densities, following the approach of Ziegler, Biersack, and Littmark (ZBL) in Ref. 13. We found that the ZBL potentials and the reproduced OF-DFT calculations have considerable differences to our more accurate self-consistent quantum chemical calculations: both the ZBL universal potential and the ZBL pair-specific calculations differ from the quantum chemical approaches by $\sim$ 5-10% above 100 eV.
Having verified the accuracy of the DMol scheme, we formed pair-specific analytical fits of the DMol screening functions to a sum of three exponentials, yielding the purely repulsive NLH potentials of this work. By comparison to the MP2 reference data, we showed that the pair-specific NLH potentials agree with the best available quantum chemical data within $\sim 2$% above 30 eV. The exponents and fitting coefficients defining the NLH potential are included in the Supplementary Material [92] and in an Open data set.
@@
The ZBL pair-specific, DMol and MP2 data sets and the NLH potential coefficients are available open access in the online Supplementary materials of this publication [92] in the compressed archive package file nlh_potentials_opendata.tar.gz , as well as at https://zenodo.org/records/14172633 (doi:0.5281/zenodo.14172632).
The online supplementary material also includes the file nlh_table_all.pdf that has a table of all the NLH potential coefficients [92].
@@ Appendix A Combining high-energy and equilibrium potentials
We now address the question of how the two types of interactions—near-equilibrium many-body and high-energy repulsive—can be combined. The importance of doing this in a systematic manner is emphasized by the conclusion of the main text that ranges in channeling conditions are sometimes sensitive to the equilibrium part of the potential (cf. section VI.1). To address this issue it is important to be able to join the repulsive and equilibrium potentials without free parameters, as the choice of such parameters could introduce additional uncertainty into the channeling results.
The key to this combination is that at the highest energies the atom collisions can practically always be handled as two-body interactions, as strong keV energy collisions involve internuclear distances $<0.5$, while regular internuclear distances are of the order of 2–3 Å. It is thus practically impossible for more than two atoms to simultaneously experience the strongly repulsive part of the potential. It is also fully sufficient to use BCA type repulsive pair potentials of the form of eq. 4 for small internuclear distances in MD simulations.
@@
where $0\leq F(r)\leq 1$ is some continuous function that has the properties $F(r)\to 1$ when $r\to 0$ and $F(r)\to 0$ when $r\to\infty$. To achieve a smooth transition from the many-body to the repulsive potential that does not affect the equilibrium potential, the function $F(r)$ should transition from 1 to 0 in a fairly narrow interval below $r_{\rm equi}$; the Fermi function is one possibility [105]. This approach has the advantage that it naturally ensures that e.g. angular dependencies associated with chemical bonds are not active at small interatomic separations.
@@
We also note that the approach to first fix the repulsive potential $V_{\rm rep}$, then construct a many-body potential by fitting the reduced energies $E_{\rm DFT}^{\rm reduced}$ could also be used to remove the ambiguity in potential joining when constructing analytical interatomic potentials.
@@ Appendix B Computation of the orbital-free density functional potentials
The basic input to the calculations are the electron densities $\rho(r)$, where $\rho$ is given in units of electrons per Å3 and the distance from the nucleus $r$ is in units of Å. We used the spherically averaged solid-state electron densities as listed in the ZBL book [13] to determine what we call the pair-specific ZBL potentials for all non-radioactive elements, plus Bi and U. In addition, we calculated electron densities for many atoms in the gas phase with the Grasp2018 program [44, 45] for comparison.
@@
Equations 24 and 25 describe the interaction of a charge cloud with a point charge, integrating over all of space, while eq. 26 describes the interaction of two charge clouds with each other. In eqs. 24 and 25, $r_{1}$ and $r_{2}$ denote the distances of the infinitesimal volume element ${\rm d}^{3}r$ from nuclei 1 and 2, respectively. Equation 26 is for now written in a general form where ${\bf x}_{1}$ and ${\bf x}_{2}$ are coordinates of two volume elements, and $|{\bf x}_{1}-{\bf x}_{2}|$ is the distance between them. The numerical value of $e^{2}/(4\pi\epsilon_{0})$ was given in the main text after eq. 1.
@@
The Coulomb integrals of eqs. 24, 25 and 26 are facile to evaluate using Green’s theorem or the Laplace expansion for the potential [49]: a spherical charge distribution yields a potential which coincides with that of a point charge
@@
Since we approximate the electron densities $\rho_{i}(r)$ by piecewise polynomials, the integrals in eqs. 27 and 40 can be carried out exactly, which improves the accuracy of the Coulomb interaction term eq. 23. This also speeds up the computation of $V_{c}$, because the integrals can be represented by another piecewise polynomial, which has to be defined only once for each atom species.
The integrals in eqs. 39, 41 and 42 were evaluated using the quad function of the SciPy library [112], using a relative tolerance of $10^{-6}$ for the quantum-mechanical terms and $10^{-7}$ for the Coulomb term $V_{ee}$. For the quantum-mechanical terms, the tolerance cannot always be reached, but the relative error usually is better than $10^{-4}$. In addition, the error estimate is not guaranteed to be accurate. Numerical problems are detected and fixed in the course of choosing the abscissae for the function $V(R)$ as described in the following.